find three consecutive positive even integers such that the product of the second and third integers is twenty more than ten times the first integer
Answer by mananth(16949)

let the first integer be x
second = x+2
third = x+4
..
(x+2)(x+4)=10x+20
x^2+6x+8=10x+20
x^2-4x-12=0
x^2-6x+4x-12=0
x(x-6)+4(x-3)=0
(x-6)(x+4)=0
x= 6
the numbers are 6,8,10.